A survey showed that 71 % of adults need correction (eyeglasses, contacts, surgery, etc.) for their eyesight. If 9 adults are randomly selected, find the probability that no more than 1 of them need correction for their eyesight. Is 1 a significantly low number of adults requiring eyesight correction?

The probability that no more than 1 of the 9 adults require eyesight correction is ________. (Round to three decimal places as needed.)

Thanks in advance!
Answer #1

Binomial Distribution

n = 9

p =0.71

q = 1 - = 0.29

P(X1) = P(X=0) + P(X=1)

So,

P(X1) =0.0003342

          = 0.000 (Round to 3 decimal places)

So,

Answer is:

0.000